The formula of a hydrate of an alkali metal carbonate may be determined by titration with standard hydrochloric acid. If 2.316g of a hydrate of sodium carbonate (Na2CO3●xH2O) is required to completely neutralize 38.49mL of hydrochloric acid (0.4198mol/L), determine the formula of the hydrate.
1
Expert's answer
2020-05-28T10:32:08-0400

Solution:

The chemical equation:

Na2CO3●xH2O(aq) + 2HCl(aq) = 2NaCl(aq) + (x+1)H2O(l) + CO2(g)

According to the chemical equation: n(Na2CO3●xH2O) = n(HCl)/2


Moles of HCl = n(HCl) = Molarity of HCl × Volume of HCl

n(HCl) = (0.4198 mol/L) × (0.03849 L) = 0.0162 mol


n(Na2CO3●xH2O) = n(HCl)/2 = (0.0162 mol) / 2 = 0.0081 mol

Molar mass of Na2CO3●xH2O = Mass of hydrate / Moles of hydrate

M(Na2CO3●xH2O) = (2.316 g) / (0.0081 mol) = 285.93 g/mol


M(Na2CO3●xH2O) = M(Na2CO3) + x●M(H2O)

The molar mass of Na2CO3 is 106 g/mol

The molar mass of H2O is 18 g/mol

x = [M(Na2CO3●xH2O) - M(Na2CO3)] / M(H2O)

x = (285.93 g/mol - 106 g/mol) / (18 g/mol) = (179.93 g//mol) / (18 g/mol) = 9.996 = 10

x = 10

Therefore, the formula of the hydrate is Na2CO3●10H2O


Answer: The formula of the hydrate is Na2CO3●10H2O